# Optical pumping of alkali-metal vapor with hyperfine-resolved buffer gas pressure

### 06 Mar 2026 v1

@@ III.2.1 Depopulation
Figure 2: Detuning-dependent factor $Q_{F}$ for $F=a$ and $F=b$ [see Eq. (24)]. Solid, dashed, and dotted curves represent $\Gamma_{\mathrm{brd}}=3,~10~\mathrm{and}~20\mathrm{~GHz}$. They converge to a common intersection at $\Delta_{a}=\Delta_{b}$, yielding $Q_{a}=Q_{b}=1$. The thick black dashed curve shows the maximum attainable $Q_{b}$ for varying $\Gamma_{\mathrm{brd}}$.
Based on the QHP-specific corrections of $\delta H^{\{\mathrm{g}\}}$ in Eq. (17), it is proved that in Liouville space, $\mathcal{A}_{\mathrm{dp}}$ similarly takes a form that links to the HP counterpart $\mathcal{A}^{(\mathrm{HP})}_{\mathrm{dp}}$
$\mathrm{Re}\Big[\mathcal{A}_{\mathrm{dp}}\Big]_{m,n}$ $=\mathrm{Re}\Big[\mathcal{A}^{(\mathrm{HP})}_{\mathrm{dp}}\Big]_{m,n}\mathcal{Q}^{\{n\}}_{\mathcal{L}},$ (25)
$\mathrm{Im}\Big[\mathcal{A}_{\mathrm{dp}}\Big]_{m,n}$ $=\mathrm{Im}\Big[\mathcal{A}^{(\mathrm{HP})}_{\mathrm{dp}}\Big]_{m,n}\mathcal{K}^{\{n\}}_{\mathcal{L}},$ (26)
where $\mathcal{X}]_{m,n}$ labels the matrix block of a superoperator $\mathcal{X}$, responsible for coupling $n$th-order to $m$th-order Zeeman coherences. The block decomposition in Liouville space serves as an efficient method for treating master equations, detailed in the Appendix. A. Here in Eqs. (25) and (26), the QHP-specific matrices $\mathcal{Q}^{\{n\}}_{\mathcal{L}}$ and $\mathcal{K}^{\{n\}}_{\mathcal{L}}$ are
$\mathcal{Q}^{\{k\}}_{\mathcal{L}}$ $=\sum_{F}\sum_{m_{F}}Q_{F}\mathcal{P}^{\{k\}}_{F}$, (27)
$\mathcal{K}^{\{k\}}_{\mathcal{L}}$ $=\sum_{F}\sum_{m_{F}}K_{F}\mathcal{P}^{\{k\}}_{F}$. (28)
This result can be mathematically viewed as a linear column transformation acting on the HP superoperator in Zeeman subspace. Together with the next section, we present that $A_{\mathrm{dp}}$ and $A_{\mathrm{rp}}$ both adopt the analogous form. The detailed proofs are provided in the Appendix. B.
@@ III.2.2 Repopulation
The repopulation term in Eq. (13) represents the generation and relaxation of the atomic ES spin during the optical pumping process. In the same way of dealing with $\mathcal{A}_{\mathrm{dp}}$, we primarily simplify $\mathcal{A}^{\{\mathrm{eg}\}}_{\mathrm{p}}$ in Eq. (13) as
$\Big[\mathcal{A}_{\mathrm{p}}^{\{\mathrm{eg}\}}\Big]_{\bar{m},n}=\Big[2(\hat{\mathbf{e}}^{*}\cdot\bm{\Delta}^{\mathrm{T}})\otimes(\bm{\Delta}^{\dagger}\cdot\hat{\mathbf{e}})\Big]_{\bar{m},n}\mathcal{Q}^{\{n\}}_{\mathcal{L}},$ (29)
Here $\bar{m}$ denotes the ES $\bar{m}$th-order Zeeman coherence subspace and its projection operator remains the same with Eq. (77) for D1 line. This result physically means that detunings $\Delta_{F}$ for F=a,bF=a,\,b contribute differently to the excitation transfer.
@@
Under lower pressure (e.g., $0.01\sim 0.1~\mathrm{amg}$ for $\mathrm{N_{2}}$), a systematic investigation in [17] offers a scale of correction from $H^{\{\mathrm{e}\}\copyright}$. For theoretical simplicity, we still adopt the approximation of Eq. (30) in the following analysis.
As already researched [1, 9], J-Damping collsion and quenching fully transfer nuclear polarization with no transfer of electronic polarization. Furthermore, any mechanism originating from the electric-dipole interaction only manipulate the electronic part of a density matrix, and assuere nuclear spin invariant, due to its short interaction timescale. So superoperators $\mathcal{A}_{\mathrm{dp}}$ and $\mathcal{A}_{\mathrm{rp}}$ share a common nuclear part
@@
Figure 3: Heat-map plot of the superoperator $\mathrm{Re}\left[A_{\mathrm{op}}\right]$ with nuclear spin $I=3/2$ and $\mathbf{s}=\hat{z}$. The underlying matrix elements are arranged with basis labeled by $|F,\,m_{F};F,\,m^{\prime}_{F})$ and presented in the HP limit. They are separated by dashed lines into $m$th- to $n$th- order Zeeman coherence blocks (see Appendix. A). The block marked by the solid lines is $\left[A_{\mathrm{op}}\right]_{0,0}$ which governs population pumping, while $\mathrm{Re}\left[A_{\mathrm{op}}\right]_{1,1}$ marked by the dotted lines contributes to evolution of the 1st-order Zeeman coherence. The vertical gray and green bands encode the detuning-dependent $Q_{F}$ for F=a,bF=a,\,b, acting in the QHP regime. Other blocks follow the same way.
Through Eqs. (10), (25), (26),and (32), the full optical pumping superoperator $\mathcal{A}_{\mathrm{op}}$ in the Zeeman subspace is
$\mathrm{Re}\Big[\mathcal{A}_{\mathrm{op}}\Big]_{{m,n}}=$ $\mathrm{Re}\Big[\mathcal{A}^{(\mathrm{HP})}_{\mathrm{op}}\Big]_{{m,n}}\mathcal{Q}^{\{n\}}_{\mathcal{L}},$ (33)
$\mathrm{Im}\Big[\mathcal{A}_{\mathrm{op}}\Big]_{{m,n}}=$ $\mathrm{Im}\Big[\mathcal{A}^{(\mathrm{HP})}_{\mathrm{op}}\Big]_{{m,n}}\mathcal{K}^{\{n\}}_{\mathcal{L}},$ (34)
where detuning-independent $\mathcal{A}^{(\mathrm{HP})}_{\mathrm{op}}$ is optical pumping superoperator in the HP limit, and its real and imaginary parts are
$\mathrm{Re}\Big[\mathcal{A}^{(\mathrm{HP})}_{\mathrm{op}}\Big]_{m,n}$ $=\left[\mathcal{A}_{\mathrm{SD}}\right]_{m,n}-\frac{\mathbf{s}}{2}\cdot\left[\bm{\mathcal{A}}_{\mathrm{SE}}\right]_{m,n},$ (35)
$\mathrm{Im}\Big[\mathcal{A}^{(\mathrm{HP})}_{\mathrm{op}}\Big]_{m,n}$ $=-i\big[\mathbf{s}\cdot\mathbf{S}^{\copyright}\big]_{m,n}.$ (36)
respectively, leading to incoherent evolution and frequency shift. In Eqs. (33) and (34), diagonal $\mathcal{Q}^{\{n\}}_{\mathcal{L}}$ and $\mathcal{K}^{\{n\}}_{\mathcal{L}}$ carry all dependence of hyperfine-resolved detuning $\Delta_{F}$. Matrix structure and detuning dependence of $\mathrm{Re}\left[\mathcal{A}_{\mathrm{op}}\right]_{m,n}$ are illustrated in Fig. 3. As for D2 line, we should make the replacement $\mathbf{s}\rightarrow-{\mathbf{s}}/{2}$ [12].
Thus, in the QHP regime, we derive a compact form of $\mathcal{A}_{\mathrm{op}}$ in Zeeman in Eq. (6), with other relaxation superoperators fomally unchanged.
III.3 Evolution in longitudinal pumping
Longitudinal pumping is an usual configuration in vapor-cell systems, where the average photon spin $\mathbf{s}$ and the static magnetic field $\mathbf{B}$ are both along the quantization axis. So we set $\mathbf{s}=s\hat{z}$ and $\mathbf{B}=B\hat{z}$ for simplicity. For atomic magnetometers, in which a weak transverse radio-frequency (RF) field is applied, evolution and steady state of $|\rho^{\{\mathrm{gg}\}})$ can be reduced to that in subspaces of the population and 1st-order Zeeman coherences. Accordingly, we concentrate on observables that manifest this steady-state behavior in experiments, including the photon absorption cross section, spin polarization, and magnetic resonance linewidth.
@@ III.3.1 Light absorption of atoms
Light absorption of atoms initially enables optical pumping, quantified by the measurable photon cross section $\sigma$. It is generally given by an induced electric dipole moment $\langle\mathbf{d}\rangle$ that oscillates with the light frequency $\omega$ and dissipates the light intensity [9]
$\sigma_{\mathrm{tot}}$ $=\dfrac{1}{\Phi\omega^{\{\mathrm{eg}\}}}\left(\Omega^{*}_{\mathrm{I}}\,\hat{\mathbf{e}}^{*}\cdot\dfrac{d}{dt}\langle\mathbf{d}\rangle+\mathrm{c.c.}\right).$ (37)
In buffer-gas cells where the ES is well adiabatically eliminated, the GS population $\left|\rho_{0}\right)=\mathcal{P}^{\{0\}}_{\mathrm{Z}}\left|\rho^{\{\mathrm{gg}\}}\right)$ that satisfies normalization $\left.\left(1^{\{\mathrm{g}\}}\right|\rho_{0}\right)=1$ can sufficiently determine $\sigma$
$\sigma$ $=-\dfrac{2}{\Phi}\left(\mathrm{Im}\left.\left[\delta H^{\{\mathrm{g}\}}\right]\right|\rho_{0}\right).$ (38)
Using $\delta H^{\{\mathrm{g}\}}$ given in Eq. (21), we have
$\sigma=\sigma_{0}(1^{\{\mathrm{g}\}}|\mathcal{Q}_{\mathcal{H}}|\rho_{0})-2{s}\sigma_{0}\left({S_{z}}\right|\mathcal{Q}_{\mathcal{H}}|\rho_{0}).$ (39)
Here $\sigma_{0}$ can be obtained as Eq. (19) by substituting $|\rho_{0})=|1^{\{\mathrm{g}\}})/(2[I])$ in Eq. (38). For arbitrary $|\rho_{0})$, Eq. (39) naturally separates $\sigma$ into two contributions. The first term represents absorption of the unpolarized hyperfine multiplets F=a,bF=a,\,b, weighted with the factors $Q_{F}$ contained in $\mathcal{Q}_{\mathcal{H}}$. The second term accounts for absorption correction of polarized atoms. In particular, population that accumulates in the stretched states $|F=a,\,m_{F}=a\rangle$ weakens the absorption of circularly polarized photons. With $s=\pm 1,\,0$ and the corresponding steady state of $|\rho_{0})$ in the section below, $\sigma_{\mathrm{lin}}$ and $\sigma_{\mathrm{cir}}$, the cross section of circular and linear polarized photons, will be obtained, respectively. They are both expressed as linear combinations of the components associated with the two multiplets
$\sigma$ $=C_{a}\,\frac{A_{0}\Gamma_{\mathrm{brd}}}{\Delta_{a}^{2}+\Gamma_{\mathrm{brd}}^{2}}+C_{b}\,\frac{A_{0}\Gamma_{\mathrm{brd}}}{\Delta_{b}^{2}+\Gamma_{\mathrm{brd}}^{2}},$ (40)
where the coefficient $C_{F}$ for F=a,bF=a,\,b is
$C_{F}$ $=p_{F}-{2s\langle S_{z}\rangle_{F}},$ (41)
with $p_{F}$ being total population of hyperfine F-multiplet and $\langle S_{z}\rangle_{F}$ describing contribution of the F-multiplet to $\langle S_{z}\rangle$
$p_{F}=$ $\sum_{m_{F}}(Fm_{F};Fm_{F}|\rho_{0}),$ (42)
$\langle S_{z}\rangle_{F}=$ $\sum_{m_{F}}(S_{z}|Fm_{F};Fm_{F})(Fm_{F};Fm_{F}|\rho_{0}).$ (43)
This means that calculations of $\sigma_{\mathrm{lin}}$ and $\sigma_{\mathrm{cir}}$ generally require a definite form of $|\rho_{0})$, which is discussed below.
III.3.2 Steady state population
Figure 4: (a) The 1-norm distance between the steady-state population and the STD with the same polarization. The curves correspond to different detuning-dependent factors $Q_{F}$ for F=a,bF=a,\,b. (b) Electron-spin polarization $P$ versus $\Gamma_{\mathrm{ex}}/\Gamma_{\mathrm{SD}}$, with color coding identical to (a) setting. (c) Steady-state population obtained with $\Gamma_{\mathrm{ex}}/\Gamma_{\mathrm{SD}}=10$, corresponding to the point marked on the red curve in (b). STD with the same polarization is shown and compared. (d) Same as (c) but corresponding to the blue marked point in (b). All results are computed for $I=3/2$ and $s=1$, fixing $R_{\mathrm{op}}/\Gamma_{\mathrm{SD}}=1$.
@@ III.3.2 Steady state population
In the presence of the light absorption and other relaxation mechanisms in Eq. (6), the population dynamics are governed by

### 14 Apr 2026 v2

@@ III.2.1 Depopulation
Figure 2: Weighting factor $Q_{F}$ for $F=a$ and $F=b$ [see Eq. (23)], with the GS hyperfine splitting $6.8~\mathrm{GHz}$ and different values of collisional broadening $\Gamma_{\mathrm{brd}}=3,~10~\mathrm{and}~20\mathrm{~GHz}$. The curves converge to a common intersection at $\Delta_{a}=\Delta_{b}$, yielding $Q_{a}=Q_{b}=1$. The thick black dashed curve shows the maximum attainable $Q_{b}$ for increasing $\Gamma_{\mathrm{brd}}$, recovering the HP limit.
The corrected depopulation superoperator in Eq. (22) reveals the physical effect arising from the difference in absorption between hyperfine split levels. Indeed, this difference is characterized by the factor $Q_{F}$ with given $F=a$ or $F=b$, as shown in Fig. 2. For hyperfine coherence between levels $F=a$ and $F=b$, e.g., $|a,m_{a};b,m_{b})$, a simple form of $\mathcal{A}_{\mathrm{dp}}$ like Eq. (22) is not available, as proved in Appendix A. Fortunately, the hyperfine coherence is usually neglected in analyzing most of the observables like spin polarization, magnetic resonance linewidth, etc, as shown in Section IV.
@@ III.2.2 Repopulation
The repopulation term in Eq. (13) represents the generation and relaxation of the atomic ES spin during the optical pumping process. In the similar way of dealing with $\mathcal{A}_{\mathrm{dp}}$, the optical excitation superoperator $\mathcal{A}^{\{\mathrm{eg}\}}_{\mathrm{p}}$ is simplified as (see Appendix A for detailed proof)
$\mathcal{A}^{\{\mathrm{eg}\}}_{\mathrm{p}}\mathcal{P}_{\mathrm{Z}}=2\sum_{F}(\mathbf{\hat{e}}^{*}\cdot\bm{d})\otimes(\bm{d}^{\dagger}\cdot\mathbf{\hat{e}})Q_{F}\mathcal{P}_{F}.$ (27)
This result physically means that detunings $\Delta_{F}$ for F=a,bF=a,\,b contribute differently to the optical excitation process.
@@
Under lower pressure (e.g., $0.01\sim 0.1~\mathrm{amg}$ for $\mathrm{N_{2}}$), a systematic investigation in [17] offers a scale of correction from $H^{\{\mathrm{e}\}\copyright}$. For theoretical simplicity, we still adopt the approximation of Eq. (28) in the following analysis. As already researched [1, 9], J-Damping collsion and quenching fully transfer nuclear polarization with no transfer of electronic polarization. Furthermore, any mechanism originating from the electric-dipole interaction only manipulate the electronic part of a density matrix, and assuere nuclear spin invariant, due to its short interaction timescale. So superoperators $\mathcal{A}_{\mathrm{dp}}$ and $\mathcal{A}_{\mathrm{rp}}$ share a common nuclear part
@@
Figure 3: Real part of optical pumping superoperator within the Zeemnan subspace, for nuclear spin $I=3/2$, D line constant $K=1$, and mean photon spin $\mathbf{s}=\hat{z}$. The underlying matrix elements are presented in the HP limit and arranged in Liouville basis $|F,\,m_{F};F,\,m^{\prime}_{F})$. They are separated by dashed lines into $m$th- to $n$th- order Zeeman coherence blocks $\mathrm{Re}\left[\mathcal{A}_{\mathrm{op}}\right]_{m,n}$ (see Appendix. B). The block marked by the solid lines ($m=n=0$) governs population dynamics, while the block marked by the dotted lines ($m=n=1$) contributes to evolution of the 1st-order Zeeman coherence. The vertical gray and green bands encode $Q_{a}$ and $Q_{b}$ Eq. (23), acting as the QHP correction. The remaining blocks follow the same structure.
Through Eqs. (10), (22), and (30), we conclude that the full optical pumping superoperator $\mathcal{A}_{\mathrm{op}}$ with the QHP correction is
$\mathcal{A}_{\mathrm{op}}\mathcal{P}_{\mathrm{Z}}=$ $\sum_{F}\left[\mathcal{A}_{\mathrm{SD}}-\frac{K\mathbf{s}}{2}\cdot\bm{\mathcal{A}}_{\mathrm{SE}}-\frac{i\Delta_{F}}{\Gamma_{\mathrm{brd}}}K\mathbf{s}\cdot\mathbf{S}^{\copyright}\right]Q_{F}\mathcal{P}_{F}.$ (31)
The compact form containing the $F$-resolved counterpart enables us to revisit the master equation Eq. (6) in the QHP regime. For higher computational efficiency and clearer physical interpretation in Liouville space, matrix elements of the superoperators are grouped into blocks according to the characteristic evolution frequencies of the density-matrix elements. The technical details of this procedure are provided in the Appendix B. Figure 3 illustrates the matrix elements and the corresponding block structure of $\mathrm{Re}\left[\mathcal{P}_{\mathrm{Z}}\mathcal{A}_{\mathrm{op}}\mathcal{P}_{\mathrm{Z}}\right]$, which represents the QHP-corrected optical pumping superoperator restricted to the Zeeman subspace of primary interest.
III.3 Observables in quasi-high-pressure regime
In this subsection, we investigate observables in the QHP regime for the common configuration in which the applied magnetic field is collinear with the pump beam. Specifically, we analyze the optical absorption, population distributions, spin polarization, and magnetic-resonance linewidth, and we highlight the ways in which their behavior departs from that observed in the HP regime.
@@ III.3.1 Light absorption of atoms
Optical absorption of atoms is quantified by the absorption cross section $\sigma$, which is expressed in terms of the effective Hamiltonian $\delta H^{\{\mathrm{g}\}}$ and the GS population $|\rho_{0})$ [9].
$\sigma$ $=-\dfrac{2}{\Phi}\left(\mathrm{Im}\left.\left[\delta H^{\{\mathrm{g}\}}\right]\right|\rho_{0}\right),$ (32)
where $(A|B)=\mathrm{Tr}[AB]$ denotes the inner product in Liouville space. Following Eqs. (20), (21) and (26), the absorption cross-section in the QHP regime is
$\sigma$ $=A\left[C_{a}\mathcal{L}(\Delta_{a})+C_{b}\mathcal{L}(\Delta_{b})\right].$ (33)
with the weighting factor $C_{F}$ defined by
$C_{F}$ $=p_{F}-{2Ks\langle S_{z}\rangle_{F}},$ (34)
where $s$ is effective photon spin along the $\hat{z}-$axis ($s=0$ or $\pm 1$ stands for the linearly or circularly polarized beam), and $p_{F}$ and $\langle S_{z}\rangle_{F}$ are the total population and the spin polarization of the states with a given $F$, i.e.
$p_{F}=$ $\sum_{m_{F}}\langle F,m_{F}|\rho_{0}|F,m_{F}\rangle,$ (35)
$\langle S_{z}\rangle_{F}=$ $\sum_{m_{F}}\langle F,m_{F}|S_{z}|F,m_{F}\rangle\langle F,m_{F}|\rho_{0}|F,m_{F}\rangle.$ (36)
The $C_{F}$ coefficient Eq. (34) comprises two terms representing distinct optical pumping processes. The $F$-resolved population term, $p_{F}$, characterizes the population redistribution between the $F=a$ and $F=b$ levels. Meanwhile, the $\langle S_{z}\rangle_{F}$ term accounts for population transfer among the $m_{F}$ sublevels, reflecting the generation of spin polarization through angular momentum exchange with incident photons. In the following of this paper, these two mechanisms are referred to as the $F$-pumping process and $m_{F}$-pumping process, respectively.
In the HP limit, the two Lorentzian lineshapes exhibit degeneracy, such that $\mathcal{L}_{a}(\Delta_{a})\approx\mathcal{L}_{b}(\Delta_{b})\approx\mathcal{L}(\Delta_{0})$, thereby suppressing the $F$-pumping process. Consequently, the $C_{F}$ coefficient reduces to the form $C=1-2Ks\langle S_{z}\rangle$, indicating that absorption is predominantly governed by spin polarization driven by $m_{F}$-pumping. In contrast, within the QHP regime, both $F$-pumping and $m_{F}$-pumping processes are integral to the system dynamics. Therefore, the $F$-resolved population, $p_{F}$, and the spin polarization, $\langle S_{z}\rangle_{F}$, encapsulate the fundamental mechanisms governing the QHP regime. Elucidating the dynamics of these parameters as a function of pump power and light frequency is crucial for uncovering the underlying principles of optical pumping.
III.3.2 Population and polarization
@@
In the presence of longitudinal pumping light, spin-destruction and spin-exchange relaxation in Eq. (6), the population dynamics are governed by